40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 61 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  14272   Tue Nov 6 09:45:32 2018 aaronMetaphysicsTreasureZojirushi is dead

New all organic machine.

  14271   Mon Nov 5 15:55:39 2018 SteveMetaphysicsTreasureZojirushi is dead

We have no coffee machine.

We are dreaming about it

We still do not have it.

Attachment 1: zoji.JPG
zoji.JPG
  14270   Mon Nov 5 13:52:18 2018 aaronUpdateComputer Scripts / Programsarm loss measuremenents

After running this script Friday night, i noticed Saturday that the data hadn't saved. Scrolling up inthe terminal, I couldn't see where I'd run the script, so I thought I'd forgotten to run it as I was making last minute changes to the scope settings Friday before leaving.

Monday it turns out I hadn't forgotten to run the script, but the script itself was getting hung up as it waited for ASS to settle, due to the offset on the ETM PIT or YAW setpoints. The script was waiting until both pitch and yaw settled to below 0.7, but yaw was reading ~15; I think this is normal, and it looks like Yuki had solved this problem by waiting for the DEMOD-OFFSET to become small, rather than just the DEMOD signal to be small. Since this is a solved problem, I think I might be using an old script, but I'm pretty sure I'm running the one in Johannes' folder that Yuki is referencing for example here. The scripts in /yutaro_scripts/ have this DEMOD-OFFSET functionality commented out, and anyway those scripts seem to do the 2D loss maps rather than 1D loss measurements.

In the meantime I blocked the beams and ran the script in DARK mode. The script is saving data in /armloss/data/run_20181105/, and runs with no exceptions thrown.

However, when I try to dither align the YARM, I get an error that "this is not a degree of freedom that has an ASS". I'm alsogetting some exceptions from MEDM about unavailable channels. It must have been something about donatella not initializing, because it's working on pianosa. I turned on YARM ADS from pianosa. Monitoring from dataviewer, I see that LSC-TRY_OUT has some spikes to 0.5, but it's mostly staying near 0. I tried returning to the previous frozen outputs, and also stepping around ETMY-[PIT/YAW] from the IFO_ALIGN screen, but didn't see much change in the behavior of LSC-TRY. I missed the other controls Gautam was using to lock before, and I've also made myself unclear on whether ASS is acting only on angular dof, or also on length.

I unblocked the beams after the DARK run was done.

  14269   Fri Nov 2 19:25:16 2018 gautamUpdateComputer Scripts / Programsloss measurements

Some facts which should be considered when doing this measurement and the associated uncertainty:

  1. When Johannes did the measurement, there was no light from the AS port diverted to the OMC. This represents ~70% loss in the absolute amount of power available for this measurement. I estimate ~1W*Tprm * Ritm * Tbs * Rbs * Tsrm * OMCsplit ~ 300uW which should still be plenty, but the real parameter of interest is the difference in reflected power between locked/no cavity situations, and how that compares to the RMS of the scope readout. For comparison, the POX DC light level is expected to be ~20uW, assuming a 600ppm AR coating on the ITMs.
  2. Even though the reflection from the arm not being measured may look like it's completely misaligned looking at the AS camera, the PDA520 which is used at the AS port has a large active area and so one must check on the oscilloscope that the other arm is truly misaligned and not hitting the photodiode to avoid interference effects artifically bloating the uncertainty.
  3. The PDA255 monitoring the MC transmission has a tiny active area. I'm not sure the beam has been centered on it anytime recently. If the beam is not well centered on that PD, and you normalize the measurements by "MC Transmission", you're likely to end up with larger error.
Quote:

This result has about 40% of uncertaintities in XARM and 33% in YARM (so big... no).

  14268   Fri Nov 2 16:42:31 2018 aaronUpdateComputer Scripts / Programsarm loss measuremenents

I'm continuing the arm loss measurements Yuki was making. I'm first familiarizing myself with the procedures for the measurement Johannes describes.

I'm not very familiar with the medm screens, so I'm just kind of poking around and checking with Gautam. I do the following:

  1. Turned Xarm ASS dither on, then off.
  2. Turned X and Y ALS on, then off shortly after
    1. Realizing I needed some guidance, I found this page on lock acquisition on the wiki
    2. Gautam showed me how to align/lock the IFO so I could take some notes, and we locked the Y arm, misaligned X.
  3. I put the PD back in the AS beam path to get the ASDC signal, and approximately centered the beam. This PD is on channel 1 of the scope, which is at 192.168.113.24.
  4. I centered the beam onto the MC2 PD that Yuki had installed. This PD is on channel 2 of the scope.
    1. Both scope channels are set to 1V scale (I also had tried 500mV, and it didn't seem to make a difference) and 10s time axis spacing (maximum integration time, since we're looking for a DC effect. Is this what we want?)
    2. The impedance for both channels is 1MOhm.
  5. I ran the script to start the loss measurement on the Y arm.
    1. python2 armloss_dcrefl_asdcpd_scope.py 192.168.113.24 1 2 5 YARM
    2. I'm reading ~15 (au?) for the MC channel and ~5% of that out the AS, which seems to make sense to me and looked to be about what Yuki the ratios when I checked the log files. However, I'm a bit confused by the normalization, because the maximum output of the MC PD is 10V, and indeed the scope's display is reading under 10V.

I've left the script running.

  14267   Fri Nov 2 12:07:16 2018 ranaUpdateCDSNDScope

https://alog.ligo-wa.caltech.edu/aLOG/index.php?callRep=44971

Let's install Jamie's new Data Viewer

  14266   Fri Nov 2 10:24:20 2018 SteveUpdatePEMroof cleaning

Physical plan is cleaning our roof and gutters today.

  14265   Fri Nov 2 09:47:57 2018 SteveMetaphysicsTreasureZojirushi is dead

     It took at least ten years to rust away. crying

Attachment 1: DSC01773.JPG
DSC01773.JPG
Attachment 2: zoji.JPG
zoji.JPG
  14264   Wed Oct 31 17:54:25 2018 gautamUpdateVACCC1 hornet power connection restored

Steve reported to me that the CC1 Hornet gauge was not reporting the IFO pressure after some cable tracing at EX. I found that the power to the unit had been accidentally disconnected. I re-connected the power and manually turned on the HV on the CC gauge (perhaps this can be automated in the new vacuum paradigm). IFO pressure of 8e-6 torr is being reported now.

Attachment 1: cc1_Hornet.png
cc1_Hornet.png
  14263   Thu Oct 25 16:17:14 2018 SteveUpdatesafetysafety training

Chub Osthelder received 40m specific basic safety traning today.

  14262   Mon Oct 22 15:19:05 2018 SteveUpdateVACMaglev controller serviced

Gautam & Steve,

Our controller is back with Osaka maintenace completed. We swapped it in this morning.

Quote:

TP-1 Osaka maglev controller  [  model TCO10M,  ser V3F04J07 ]  needs maintenance. Alarm led  on indicating  that we need Lv2 service.

The turbo and the controller are in good working order.

*****************************

Hi Steve,

Our maintenance level 2 service price is $...... It consists of a complete disassembly of the controller for internal cleaning of all ICB’s, replacement of all main board capacitors, replacement of all internal cooling units, ROM battery replacement, re-assembly, and mandatory final testing to make sure it meets our factory specifications. Turnaround time is approximately 3 weeks.

  RMA 5686 has been assigned to Caltech’s returning TC010M controller. Attached please find our RMA forms. Complete and return them to us via email, along with your PO, prior to shipping the cont

Best regards,

Pedro Gutierrez

Osaka Vacuum USA, Inc.

510-770-0100 x 109

*************************************************

our TP-1 TG390MCAB is 9 years old. What is the life expectancy of this turbo?

                        The Osaka maglev turbopumps are designed with a 100,000 hours(or ~ 10 operating years) life span but as you know most of our end-users are

                        running their Osaka maglev turbopumps in excess of 10+, 15+ years continuously.     The 100,000 hours design value is based upon the AL material being rotated at

                        the given speed.   But the design fudge factor have somehow elongated the practical life span.  

We should have the cost of new maglev & controller in next year budget. I  put the quote into the wiki.

 

                         

 

 

Attachment 1: our_controller_is_back.png
our_controller_is_back.png
  14261   Thu Oct 18 00:27:37 2018 KojiUpdateSUSSUS PD Whitening board inspection

[Gautam, Koji]

As a part of the preparation for the replacement of c1susaux with Acromag, I made inspection of the coil-osem transfer function measurements for the vertex SUSs.

The TFs showed typical f^-2 with the whitening on except for ITMY UL (Attachment 1). Gautam told me that this is a known issue for ~5 years.
We made a thorough inspection/replacement of the components and identified the mechanism of the problem.
It turned out that the inputs to MAX333s are as listed below.

  Whitening ON Whitening OFF
UL ~12V ~8.6V
LL 0V 15V
UR 0V 15V
LR 0V 15V
SD 0V 15V

The switching voltage for UL is obviously incorrect. We thought this comes from the broken BIO board and thus swapped the corresponding board. But the issue remained. There are 4 BIO boards in total on c1sus, so maybe we have replaced a wrong board?

Initially, we thought that the BIO can't drive the pull-up resistor of 5KOhm from 15V to 0V (=3mA of current). So I have replaced the pull-up resistor to be 30KOhm. But this did not help. These 30Ks are left on the board.
 

Attachment 1: 43.png
43.png
  14260   Wed Oct 17 20:46:24 2018 yukiConfigurationASCY end table upgrade plan

To do for Green Locking in YARM:

The auto-alignment servo should be completed. This servo requires many parameters to be optimized: demodulation frequency, demodulation phase, servo gain (for each M1/2 PIT/YAW), and matrix elements which can remove PIT-YAW coupling. 

  14259   Wed Oct 17 09:31:24 2018 SteveUpdatePSLmain laser off

The main laser went off when PSL doors were opened-closed. It was turned back on and the PSL is locked.

Attachment 1: Inno2wFlipped_off.png
Inno2wFlipped_off.png
  14258   Tue Oct 16 00:44:29 2018 yukiUpdateComputer Scripts / Programsloss measurements

The scripts for measuring armloss are in the directory "/opt/rtcds/caltech/c1/scripts/lossmap_scripts/armloss_scope".

  • armloss_derefl_asdcpd_scope.py: gets data and makes ascii file.
  • armloss_AS_calc.py: calculates armloss from selected a set of files.
  • armloss_calc_histogram.py: calculates armloss from selected files and makes histogram.
  14257   Mon Oct 15 20:11:56 2018 yukiConfigurationASCY end table upgrade plan

Final Procedure Report for Green Locking in YARM:

Purpose

The current setup of AUX Y-arm Green locking has to be improved because:

  • current efficiency of mode matching is about 50%
  • current setup doesn't separate the degrees of freedom of TEM01 with PZT mirrors (the difference of gouy phase between PZT mirrors should be around 90 deg) 
  • we want to remotely control PZT mirrors for alignment

What to do

  • Design the new setup and order optices needed (finished!)
      - As the new setup I designed, adding a new lens and slightly changing the position of optics are only needed. The new lens was arrived here.
  • Check electronics (PZT, PZT driver, high voltage, cable, anti-imaging board) (finished!)
      
    - All electronics were made sure performing well.
      - The left thing to do is making a cable. (Today's tasks)
  • Calibrate PZT mirror [mrad/V] (finished!)
      
    - The result was posted here --> elog:40m/14224.
  • Measure the status value of the current setup (power of transmitted light ...etc) (finished!)
  • Install them in the Y-end table and align the beam (Almost finished!) (GTRY signal is 0.3 which means Mode-Matching efficiency is about 30%. It should be improved.)
  • Measure the status value of the new setup (finished!)
  • Prepare the code of making alignment automaticaly
    • see sitemap.adl>ASC>c1asy. I prepared medm. If you move PZT SLIDERS then you can see the green beam also moves.
    • Preparing filters is needed. You can copy them from C1ASX.
    • Note that now you cannot use C1ASX servo because filters are not applied.
  14256   Mon Oct 15 13:59:42 2018 SteveUpdateVACdrypump replaced

Steve & Bob,

Bob removed the head cover from the housing to inspect the condition of the the tip seal. The tip seal was fine but the viton cover seal had a bad hump. This misaligned the tip seal and it did not allow it to rotate.

It was repositioned an carefully tithened. It worked. It's starting current transiant measured 28 A and operational mode 3.5 A

This load is normal with an old pump. See the brand new DIP7 drypump as spare was 25 A at start and  3.1 A in operational mode. It is amazing how much punishment a slow blow ceramic 10A  fuse can take [ 0215010.HXP ]

In the future one should measure the current pick up [ transient <100ms ] after the the seal change with Fluke 330 Series Current Clamp

 

It was swapped in and the foreline pressure dropped to 24 mTorr after 4 hours. It is very good. TP3 rotational drive current  0.15 A at 50K rpm   24C

Quote:

Gautam and Steve,

Our TP3 drypump seal is at 360 mT [0.25A load on small turbo]  after one year.  We tried to swap in old spare drypump with new tip  seal. It was blowing it's fuse, so we could not do it.

Noisy aux drypump turned on and opened to TP3 foreline [ two drypumps are in  the foreline now ]  The pressure is 48 mT and 0.17A load on small turbo.

 

Attachment 1: drypump_swap.png
drypump_swap.png
  14255   Mon Oct 15 12:52:54 2018 yukiUpdateComputer Scripts / Programsadditional comments
Quote:

but there's one weirdness: It get's the channel offset wrong. However this doesn't matter in our measurement because we're subtracting the dark level, which sees the same (wrong) offset.

When you do this measurement with oscilloscope, take care two things:

  1. set y-range of scope as to every signal fits in display: otherwise the data sent from scope would be saturated.
  2. set y-position of scope to the center and don't change it; otherwise some offset would be on the data.
  14254   Mon Oct 15 10:32:13 2018 yukiUpdateComputer Scripts / Programsloss measurements

I used these values for measuring armloss:

  • Transmissivitity of ITM = 1.384e-2 * (1 +/- 1e-2) 
  • Transmissivitity of ETM = 13.7e-6 * (1 +/- 5e-2)
  • Mode-Matching efficiency of XARM = 0.912 * (1 +/- 2e-2)
  • Mode-Matching efficiency of YARM = 0.867 * (1 +/- 2e-2)
  • modulation depth m1 (11MHz) = 0.179 * (1 +/- 2e-2)
  • modulation depth m2 = 0.226 * (1 +/- 2e-2),

then the uncertainties reported by the individual measurements are on the order of 6 ppm (~6.2 for the XARM, ~6.3 for the YARM). This accounts for fluctuations of the data read from the scope and uncertainties in mode-matching and modulation depths in the EOM. I made histograms for the 20 datapoints taken for each arm: the standard deviation of the spread is over 6ppm. We end up with something like:

XARM: 123 +/- 50 ppm
YARM: 152+/- 50 ppm

This result has about 40% of uncertaintities in XARM and 33% in YARM (so big... no).

In the previous measurement, the fluctuation of each power was 0.1% and the fluctuation of P(Locked)/P(misaligned) was also 0.1%. Then the uncertainty was small. On the other hand in my measurement, the fluctuation of power is about 2% and the fluctuation of P(Locked)/P(misaligned) is 2%. That's why the uncertainty became big.

We want to measure tiny value of loss (~100ppm). So the fluctuation of P(Locked)/P(misaligned) must be smaller than 1.6%.

(Edit on 10/23)
I think the error is dominated by systematic error in scope. The data of beam power had only 3 degits. If P(Locked) and P(misaligned) have 2% error, then
\frac{P_L}{P_M}\frac{1}{1+T_{\mathrm{ITM}}} = 0.99(3).
You have to check the configuration of scope.

Attachment 1: XARM_20181015_1500.pdf
XARM_20181015_1500.pdf
Attachment 2: YARM_20181015_1500.pdf
YARM_20181015_1500.pdf
  14253   Sun Oct 14 16:55:15 2018 not gautamUpdateCDSpianosa upgrade

DASWG is not what we want to use for config; we should use the K. Thorne LLO instructions, like I did for ROSSA.

Quote:

pianosa has been upgraded to SL7. I've made a controls user account, added it to sudoers, did the network config, and mounted /cvs/cds using /etc/fstab. Other capabilities are being slowly added, but it may be a while before this workstation has all the kinks ironed out. For now, I'm going to follow the instructions on this wiki to try and get the usual LSC stuff working.

  14251   Sat Oct 13 20:11:10 2018 yukiUpdateComputer Scripts / Programsloss measurements
Quote:

the script "armloss_AS_calc.py",

  • "ETM_YAW_L_DEMOD_I_OUTPUT" was little noisy even when the arm was locked.
  • The reflected beam power when locked was higher than when misaligned. It seemed strange for me at first. Johannes suggested that it was caused by over-coupling cavity. It is possible when r_{ETMY}>>r1_{ITMY}.

Some changes were made in the script for getting the signals of beam power:

  • The script sees "C1:ASS-X(Y)ARM_ETM_PIT/YAW_L_DEMOD_I_OUTPUT" and stops running until the signals become small, however some offset could be on the signal. So I changed it into waiting until (DEMOD - OFFSET) becomes small. (Yesterday I wrote ETM_YAW_L_DEMOD_I_OUTPUT was about 15 and was little noisy. I was wrong. That was just a offset value.)
  • I added a code which stops running the script when the power of transmitted IR beam is low. You can set this threshold. The nominal value of "C1:LSC-TRX(Y)_OUT16" is 1.2 (1.0), so the threshold is set 0.8 now.  

In the yesterday measurement the beam power of ASDC is higher when locked than when misaligned and I wrote it maybe caused by over-coupled cavity. Then I did a calculation as following to explain this:

  • assume power transmissivity of ITM and ETM are 1.4e-2 and 1.4e-5.
  • assume loss-less mirror, you can calculate amplitude reflectivity of ITM and ETM.
  • consider a cavity which consists two mirrors and is loss-less, then \frac{E_{r}}{E_{in}} = \frac{-r_1+r_2e^{i\phi}}{1-r_1r_2e^{i\phi}} holds. r1 and r2 are amplitude reflectivity of ITM and ETM, and E is electric filed.
  • Then you can calculate the power of reflected beam when resonated and when anti-resonated. The fraction of these value is \frac{P_{RESONANT}}{P_{ANTI-RESO}} = 0.996, which is smaller than 1.
  • I found this calculation was wrong! Above calculatation only holds when cavity is aligned, not when misaligned. 99.04% of incident beam power reflects when locked, and (100-1.4)% reflects when misaligned. The proportion is P(locked)/P(misaligned)=1.004, higher than 1.

 

  14248   Fri Oct 12 20:20:29 2018 yukiUpdateComputer Scripts / Programsloss measurements

I ran the script for measuring arm-loss and calculated rough Y-arm round trip loss temporally. The result was 89.6ppm. (The error should be considered later.)

The measurement was done as follows:

  1. install hardware
    1. Put a PD (PDA520) in anti-symmetric beam path to get ASDC signal.
    2. Use a PD (PDA255) in MC2 box to get tranmitted light of IMC. It is used to normalize the beam power.
    3. Connect those 2 PDs to oscilloscope (IP: 192.168.113.25) and insert an internet cable to it.
  2. measure DARK noise
    1. Block beam going into PDs with dampers and turn off the room light.
    2. Run the script "armloss_dcrefl_acdcpd_scope.py" using "DARK" mode.
  3. measure the ASDC power when Y-arm locked and misaligned
    1. Remove dampers and turn off the room light.
    2. Dither-align the interferometer with both arms locked. Freeze outputs when done. (Click C1ASS.adl>!MoreScripts>ON and click C1ASS.adl>!MoreScripts>FreezeOutputs.)
    3. Misalign ETMX + ITMX. (Just click "Misalign" button.)
    4. Further misalign ITMX with the slider. (see previous study: ITMX needs to be misaligned further. Moving the slider by at least +0.2 is plentiful to not have the other beam interfere with the measurement.)
    5. Start the script "armloss_dcrefl_acdcpd_scope.py" using "ETMY" mode, which does the following:
      1. Resume dithering of the YARM.
      2. Check YARM dither error signal rms with CDS. If they're calm enough, proceed. (In the previous study the rms threshold was 0.7. Now "ETM_YAW_L_DEMOD_I" signal was 15 (noisy), then the threshold was set 17.)
      3. Freeze dithering.
      4. Start a new set of averages on the scope, wait T_WAIT (5 seconds).
      5. Read data (= ASDC power and MC2 trans) from scope and save.
      6. Misalign ETMY and wait 5s. (I added a code which switchs LSC mode ON and OFF.)
      7. Read data from scope and save.
      8. Repeat desired amount of times.
  4. calculate the arm loss
    1. Start the script "armloss_AS_calc.py", whose content is follows:
      • requires given parameters: Mode-Matching effeciency, modulation depth, transmissivity. I used the same value as Johannes did last year, which are (huga)
      • reads datafile of beam power at ASDC and MC2 trans, which file is created by "armloss_dcrefl_acdcpd_scope.py".
      • calculates arm loss from the equation (see 12528 and 12854).

Result:

YARM
('AS_DARK =', '0.0019517200000000003') #dark noise at ASDC 
('MC_DARK =', '0.02792') #dark noise at MC2 trans
('AS_LOCKED =', '2.04293') #beam power at ASDC when the cavity was locked 
('MC_LOCKED =', '2.6951620000000003')
('AS_MISALIGNED =', '2.0445439999999997') #beam power at ASDC when the cavity was misaligned
('MC_MISALIGNED =', '2.665312')

\hat{P} = \frac{P_{AS}-P_{AS}^{DARK}}{P_{MC}-P_{MC}^{DARK}} #normalized beam power 

\hat{P}^{LOCKED}=0.765,\ \hat{P}^{MISALIGNED}=0.775,\ \mathcal{L}=89.6\ \mathrm{ppm}

Comments:

  • "ETM_YAW_L_DEMOD_I_OUTPUT" was little noisy even when the arm was locked.
  • The reflected beam power when locked was higher than when misaligned. It seemed strange for me at first. Johannes suggested that it was caused by over-coupling cavity. It is possible when r_{ETMY}>>r1_{ITMY}.
  • My first (wrong) measurement said the arm loss was negative(!). That was caused by lack of enough misalignment of another arm mirrors. If you don't misalign ITMX enough then the beam or scattered light from X-arm would bring bad. The calculated negative loss would be appeared only when \frac{\hat{P}^{LOCKED}}{\hat{P}^{MISALIGNED}} > 1 + T_{ITM}
  • Error should be considered.
  • Parameters given this time should be measured again. 
  14247   Fri Oct 12 17:37:03 2018 SteveUpdateVACpressure gauge choices

We want to measure the pressure gradient in the 40m IFO

Our old MKS cold cathodes are out of order. The existing working gauge at the pumpspool is InstruTech CCM501

The plan is to purchase 3 new gauges for ETMY, BS and MC2 location.

Basic cold cathode     or    Bayard-Alpert Pirani

    

 

  14245   Fri Oct 12 12:29:34 2018 yukiUpdateComputer Scripts / Programsloss measurements

With Gautam's help, Y-arm was locked. Then I ran the script "armloss_dcrefl_asdcpd_scope.py" which gets the signals from oscilloscope. It ran and got data, but I found some problems.

  1. It seemed that a process which makes arm cavity mislaigned in the script didn't work.
  2. The script "armloss_dcrefl_asdcpd_scope.py" gets the signal and the another script "armloss_AS_calc.py" calculates the arm loss. But output file the former makes doesn't match with the type the latter requires. A script converts format is needed.

Anyway, I got the data needed so I will calculate the loss after converting the format.

  14244   Fri Oct 12 08:27:05 2018 SteveUpdateVACdrypump

Gautam and Steve,

Our TP3 drypump seal is at 360 mT [0.25A load on small turbo]  after one year.  We tried to swap in old spare drypump with new tip  seal. It was blowing it's fuse, so we could not do it.

Noisy aux drypump turned on and opened to TP3 foreline [ two drypumps are in  the foreline now ]  The pressure is 48 mT and 0.17A load on small turbo.

Attachment 1: forepump.png
forepump.png
  14243   Thu Oct 11 13:40:51 2018 yukiUpdateComputer Scripts / Programsloss measurements
Quote:

This is the procedure I follow when I take these measurements for the XARM (symmetric under XARM <-> YARM):

  1. Dither-align the interferometer with both arms locked. Freeze outputs when done.
  2. Misalign ETMY + ITMY.
  3. ITMY needs to be misaligned further. Moving the slider by at least +0.2 is plentiful to not have the other beam interfere with the measurement.
  4. Start the script, which does the following:
    1. Resume dithering of the XARM
    2. Check XARM dither error signal rms with CDS. If they're calm enough, proceed.
    3. Freeze dithering
    4. Start a new set of averages on the scope, wait T_WAIT (5 seconds)
    5. Read data (= ASDC power and MC2 trans) from scope and save
    6. Misalign ETMX and wait 5s
    7. Read data from scope and save
    8. Repeat desired amount of times
  5. Close the PSL shutter and measure the PD dark levels

Information for the armloss measurement:

  • Script which gets the data:  /users/johannes/40m/armloss/scripts/armloss_scope/armloss_dcrefl_asdcpd_scope.py
  • Script which calculates the loss: /users/johannes/40m/armloss/scripts/misc/armloss_AS_calc.py
  • Before doing the procedure Johannes wrote you have to prepare as follows:
    • put a PD in anti-symmetric beam path to get ASDC signal.
    • put a PD in MC2 box to get tranmitted light of IMC. It is used to normalize the beam power.
    • connect those 2 PDs to oscilloscope and insert an internet cable to it.
  • Usage: python2 armloss_dcrefl_asdcpd_scope.py [IP address of Scope] [ScopeCH for AS] [ScopeCH for MC] [Num of iteration] [ArmMode]

Note: The scripts uses httplib2 module. You have to install it if you don't have.

The locked arms are needed to calculate armloss but the alignment of PMC is deadly bad now. So at first I will make it aligned. (Gautam aligned it and PMC is locked now.) 

gautam: The PMC alignment was fine, the problem was that the c1psl slow machine had become unresponsive, which prevented the PMC length servo from functioning correctly. I rebooted the machine and undid the alignment changes Yuki had made on the PSL table.

  14241   Wed Oct 10 12:38:27 2018 yukiConfigurationLSCAll hardware was installed

I connected DAC - AIboard - PZTdriver - PZT mirrors and made sure the PZT mirrors were moving when changing the signal from DAC. Tomorrow I will prepare alignment servo with green beam for Y-arm.

  14240   Tue Oct 9 23:03:43 2018 yukiConfigurationLSCYarm Green locking was recovered

[ Yuki, Gautam, Steve ]

To align the green beam in Y-end these hardware were installed:

  • PZT mirrors in Y-end table
  • PZT driver in 1Y4 rack
  • Anti-Imaging board in 1Y4 rack
  • cables (DAC - AIboard - PZTdriver - PZT)
  • high voltage supplier 

I made sure that DAC CH9~16 and cable to AI-board worked correctly. 

When we applied +100V to PZT driver and connected DAC, AI-board and PZT drive, the output voltage of the driver was not correct. I'll check it tomorrow.

Attachment 1: Pic_1Y4.jpg
Pic_1Y4.jpg
Attachment 2: Pic_PZTcable.jpg
Pic_PZTcable.jpg
  14239   Tue Oct 9 16:05:29 2018 gautamConfigurationASCc1tst deleted, c1asy deployed.

Setting up c1asy:

  • Backed up old c1tst.mdl as c1tst_old_bak.mdl in /opt/rtcds/userapps/release/cds/c1/models
  • Copied the c1tst model to /opt/rtcds/userapps/release/isc/c1/models/c1asy.mdl as this is where the c1asx.mdl file resides.
  • Backed up original c1rfm.mdl as c1rfm_old.mdl in /opt/rtcds/userapps/release/cds/c1/models (since the old c1tst had an RFM block which is unnecessary).
  • Deleted offending RFM block from c1rfm.mdl.
  • Recompiled and re-installed c1rfm.mdl. Model has not yet been restarted, as I'd like suspension watchdogs to be shutdown, but c1susaux EPICS channels are presently not responsive.
  • Removed c1tst model (C-node91) from /opt/rtcds/caltech/c1/target/gds/param/testpoints.
  • Removed /opt/rtcds/caltech/c1/target/gds/param/tpchn_c1tst.par (at this point, DCUID 91 is free for use by c1asy).
  • Moved c1tst line in /opt/rtcds/caltech/c1/target/daqd/master to "old model definitions models" section.
  • Added /opt/rtcds/caltech/c1/target/gds/param/tpchn_c1asy.par to the master file.
  • Edited/diskless/root.jessie/etc/rtsystab to allow c1asy to be run on c1iscey.
  • Finally, I followed the instructions here to get the channels into frames and make all the indicators green.

Now Yuki can work on copying the simulink model (copy c1asx structure) and implementing the autoalignment servo.

Attachment 1: CDSoverview_ASY.png
CDSoverview_ASY.png
  14238   Mon Oct 8 18:56:52 2018 gautamConfigurationASCc1asx filter coefficient file missing

While pointing Yuki to the c1asx servo system, I noticed that the filter file for c1asx is missing in the usual chans directory. Why? Backups for it exist in the filter_archive subdirectory. But there is no current file. Clearly this doesn't seems to affect the realtime code execution as the ASX model seems to run just fine. I copied the latest backup version from the archive area into the chans directory for now.

  14237   Mon Oct 8 00:46:35 2018 yukiConfigurationASCY-end table upgrade
Quote:

I measured it with the wrong setting of a powermeter. The correct ones are here:

  • beam power going into doubling crystal (red12): 240 mW, 1064nm
  • beam power transmitted dichroic mirror (Y5): 0.70 mW, 532nm
  • beam power going into faraday isolator (green5): 0.58 mW, 532nm
  • beam power going out from faraday isolator (green5): 0.54 mW, 532nm
  • beam power going to ETMY: 0.37 mW, 532nm
  • beam power of transmitted green light of Y-arm, which was measured by C1:ALS-TRY_OUT: 0.5 (see attachment #1)

After installation I measured these power again.

  • beam power going into doubling crystal: 241 mW, 1064nm
  • beam power transmitted dichroic mirror: 0.70 mW, 532nm
  • beam power going into faraday isolator: 0.56 mW, 532nm
  • beam power going out from faraday isolator: 0.53 mW, 532nm
  • beam power going to ETMY: 0.36 mW, 532nm

There is a little power loss. That may be because of adding one lens in the beam path. I think it is allowable margin.

  14236   Sun Oct 7 22:30:42 2018 yukiConfigurationLSCYarm Green locking was recovered

I finished installation of optics in the Y-end and recovered green locking. Current ALS-TRY_OUTPUT is about 0.25, which is lower than before. So I still continue the alignment of the beam. The simulation code was attached. (Sorry. The optic shown as QWP2 is NOT QWP. It's HWP.)

Attachment 1: Pic_NewLayout1007.jpg
Pic_NewLayout1007.jpg
Attachment 2: YendGreenModeMatching.zip
  14235   Sun Oct 7 16:51:03 2018 gautamConfigurationLSCYarm triggering changed

To facilitate Yuki's alignment of the EY green beam into the Yarm cavity, I have changed the LSC triggering and PowNorm settings to use only the reflected light from the cavity to do the locking of Arm Cavity length to PSL. Running the configure script should restore the usual TRY triggering settings. Also, the X arm optics were macroscopically misaligned in order to be able to lock in this configuration.

  14234   Fri Oct 5 22:49:22 2018 yukiConfigurationASCY-end table upgrade

I designed a new layout. It has good mode-matching efficiency, reasonable sensitivity to component positions, good Gouy phase separation. I'm setting optics in the Y-end table. The layout will be optimized again after finishing (rough) installation.  (The picture will be posted later)

  14233   Fri Oct 5 17:47:55 2018 gautamConfigurationASCY-end table upgrade

What about just copying the Xend layout? I think it has good MM (per calculations), reasonable (in)sensitivity to component positions, good Gouy phase separation, and I think it is good to have the same layout at both ends. Since the green waist has the same size and location in the doubling crystal, it should be possible to adapt the X end solution to the Yend table pretty easily I think.

Quote:

The setup I designed is here. It can bring 100% mode-matching and good separation of degrees of TEM01, however I found a probrem. The picture of setup is attached #3. You can see the reflection angle at Y7 and Y8 is not appropriate. I will consider the schematic again.

  14232   Fri Oct 5 17:32:38 2018 yukiConfigurationASCY-end table upgrade

I measured it with the wrong setting of a powermeter. The correct ones are here:

  • beam power going into doubling crystal (red12): 240 mW, 1064nm
  • beam power transmitted dichroic mirror (Y5): 0.70 mW, 532nm
  • beam power going into faraday isolator (green5): 0.58 mW, 532nm
  • beam power going out from faraday isolator (green5): 0.54 mW, 532nm
  • beam power going to ETMY: 0.37 mW, 532nm
  • beam power of transmitted green light of Y-arm, which was measured by C1:ALS-TRY_OUT: 0.5 (see attachment #1)

The calculated conversion efficiency of SHG crystal is 1.2%W.

  14231   Fri Oct 5 00:46:17 2018 KojiConfigurationASCY-end table upgrade

???

The SHG crystal has the conversion efficiency of ~2%W (i.e. if you have 1W input @1064, you get 2% conversion efficiency ->20mW@532nm)

It is not possible to produce 0.58mW@532nm from 20.9mW@1064nm because this is already 2.8% efficiency.

 

  14230   Thu Oct 4 22:15:30 2018 yukiConfigurationASCY-end table upgrade

Before changing setup at Y-end table, I measured the status value of the former setup as follows. These values will be compared to those of upgraded setup.

  • beam power going into doubling crystal (red12): 20.9 mW with filter, 1064nm
  • beam power going out from doubling crystal (red12): 26.7 mW with filter, 532nm
  • beam power going into faraday isolator (green5): 0.58 mW without filter, 532nm
  • beam power going out from faraday isolator (green5): 0.54 mW without filter, 532nm
  • beam power going to ETMY: 0.37 mW without filter, 532nm
  • beam power of transmitted green light of Y-arm, which was measured by C1:ALS-TRY_OUT: 0.5 (see attachment #1)

(These numbers are shown in the attachment #2.)

The setup I designed is here. It can bring 100% mode-matching and good separation of degrees of TEM01, however I found a probrem. The picture of setup is attached #3. You can see the reflection angle at Y7 and Y8 is not appropriate. I will consider the schematic again.

Attachment 1: AUXYGreenLock20180921.jpg
AUXYGreenLock20180921.jpg
Attachment 2: Pic_FormerSetup.jpeg
Pic_FormerSetup.jpeg
Attachment 3: Pic_CurrentSetup1004.jpg
Pic_CurrentSetup1004.jpg
  14229   Thu Oct 4 08:25:50 2018 SteveUpdateVACrga scan pd81 at day 78

 

 

Attachment 1: pd81d78.png
pd81d78.png
  14228   Thu Oct 4 00:44:50 2018 yukiConfigurationASCAI board improvement

[ Yuki, Gautam ]

I made a cable which connects DAC port (40 pins) and AI board (25 pins). I will check if it works.

Tomorrow I will change setup for improvement of AUX Y-end green locking. Any optics for IR will not be moved in my design, so this work doesn't affect Y-arm locking with main beam. 
While doing this work, I will do:

  • check if the cable works
  • make another cable which connects AI board (10 pins) and PZT driver (10 pins).
  • check if eurocate in Y-rack (IY4?) applies +/-5V, +/-15V and +/-24V. It will be done using an expansion card.
  • improve alignment servo for X-end.
  • setup alignment servo for Y-end.
  • about optical loss measurement.  
  14227   Wed Oct 3 18:15:34 2018 yukiConfigurationASCAI board improvement

[ Yuki, Gautam ]

I improved Anti-Imaging board (D000186-Rev.D), which will be put between DAC port and PZT driver board.

It had notches at f = 16.6 kHz and 32.7 kHz, you can see them in the plot attached. So I replaced some resistors as follows:

  • R6 and R7 replaced with 511 ohm (1206 thin film resistor)
  • R8 replaced with 255 ohm (1206 thin film resistor)
  • R14 and R15 replaced with 549 ohm (1206 thin film resistor)
  • R16 replaced with 274 ohm (1206 thin film resistor)

Then the notch moved to 65.9 kHz (> sampling frequency of DAC = 64 kHz, good!). 
(The plot enlarged around the notch frequency and the plot of all channels will be posted later.)

All electronics and optics seem to be ready. 

Reference, elog:40m/8857
Diagram, D000186-D.pdf

Attachment 1: TF_AIboard.pdf
TF_AIboard.pdf
  14226   Wed Oct 3 14:24:40 2018 yukiConfigurationASCY end table upgrade plan

Interim Procedure Report:

Purpose

The current setup of AUX Y-arm Green locking has to be improved because:

  • current efficiency of mode matching is about 50%
  • current setup doesn't separate the degrees of freedom of TEM01 with PZT mirrors (the difference of gouy phase between PZT mirrors should be around 90 deg) 
  • we want to remotely control PZT mirrors for alignment

What to do

  • Design the new setup and order optices needed (finished!)
      - As the new setup I designed, adding a new lens and slightly changing the position of optics are only needed. The new lens was arrived here.
  • Check electronics (PZT, PZT driver, high voltage, cable, anti-imaging board) (finished!)
      
    - All electronics were made sure performing well.
      - The left thing to do is making a cable. (Today's tasks)
  • Calibrate PZT mirror [mrad/V] (finished!)
      
    - The result was posted here --> elog:40m/14224.
  • Measure the status value of the current setup (power of transmitted light ...etc) (Tomorrow, --> finished!)
  • Install them in the Y-end table and align the beam (Will start from Tomorrow) (The setup has a probrem I found on 10/04)
  • Measure the status value of the new setup
      - I want to finish above during my stay.
  • Prepare the code of making alignment automaticaly
  14225   Tue Oct 2 23:57:16 2018 gautamUpdatePonderSqueezeSqueezing scenarios

[kevin, gautam]

We have been working on double checking the noise budget calculations. We wanted to evaluate the amount of squeezing for a few different scenarios that vary in cost and time. Here are the findings:

Squeezing scenarios

Sqz [dBvac] fmin [Hz] PPRM [W] PBS [W] TPRM [%] TSRM [%]
-0.41 215 0.8 40 5.637 9.903
-0.58 230 1.7 80 5.637 9.903
-1.05 250 1.7 150 1 17
-2.26 340 10 900 1 17

All calculations done with

  • 4.5kohm series resistance on ETMs, 15kohms on ITMs, 25kohm on slow path on all four TMs.
  • Detuning of SRC = -0.01 deg.
  • Homodyne angle = 89.5 deg.
  • Homodyne QE = 0.9. 
  • Arm losses is 20ppm RT.
  • LO beam assumed to be extracted from PR2 transmission, and is ~20ppm of circulating power in PRC.

Scenarios:

  1. Existing setup, new RC folding mirrors for PRG of ~45.
  2. Existing setup, send Innolight (Edwin) for repair (= diode replacement?) and hope we get 1.7 W on back of PRM.
  3. Repair Innolight, new PRM and SRM, former for higher PRG, latter for higher DARM pole.
  4. Same as #3, but with 10 W input power on back of PRM (i.e. assuming we get a fiber amp).

Remarks:

  • The errors on the small dB numbers is large - 1% change in model parameters (e.g. arm losses, PRG, coil driver noise etc) can mean no observable squeezing. 
  • Actually, this entire discussion is moot unless we can get the RIN of the light incident on the PRM lower than the current level (estimated from MC2 transmission, filtered by CARM pole and ARM zero) by a factor of 60dB.
    • This is because even if we have 1mW contrast defect light leaking through the OMC, the beating of this field (in the amplitude quadrature) with the 20mW LO RIN (also almost entirely in the amplitude quad) yields significant noise contribution at 100 Hz (see Attachment #1).
    • Actually, we could have much more contrast defect leakage, as we have not accounted for asymmetries like arm loss imbalance.
    • So we need an ISS that has 60dB of gain at 100 Hz. 
    • The requirement on LO RIN is consistent with Eq 12 of this paper.
  • There is probably room to optimize SRC detuning and homodyne angle for each of these scenarios - for now, we just took the optimized combo for scenario #1 for evaluating all four scenarios.
  • OMC displacement noise seems to only be at the level of 1e-22 m/rtHz, assuming that the detuning for s-pol and p-pol is ~30 kHz if we were to lock at the middle of the two resonances
    • This assumes 0.02 deg difference in amplitude reflectivity b/w polarizations per optic, other parameters taken from aLIGO OMC design numbers.
    • We took OMC displacement noise from here.

Main unbudgeted noises:

  • Scattered light.
  • Angular control noise reinjection (not sure about the RP angular dynamics for the higher power yet).
  • Shot noise due to vacuum leaking from sym port (= DC contrast defect), but we expect this to not be significant at the level of the other noises in Atm #1.
  • Osc amp / phase.
  • AUX DoF cross coupling into DARM readout.
  • Laser frequency noise (although we should be immune to this because of our homodyne angle choice).

Threat matrix has been updated.

Attachment 1: PonderSqueeze_NB_LORIN.pdf
PonderSqueeze_NB_LORIN.pdf
  14224   Tue Oct 2 18:50:53 2018 yukiConfigurationASCPZT mirror calibration

[ Yuki, Gautam ]

I calibrated PZT mirrors. The ROUGH result was attached. (Note that some errors and trivial couplings coming from inclination of QPD were not considered here. This should be revised and posted again.) 

The PZT mirrors I calibrated were:

  • A 2-inch CVI mirror (45 degree, HR and AR for 532nm)
  • A 1-inch Laseroptik mirror (45 degree, HR and AR for 532nm)

I did the calibration as follows:

  • +-15V was supplied to PZT driver circuit, +100V to PZT driver bias, and +-18V to QPD amplifier.
  • Optical path length was set to be same as that when I calibrated QPD, which is 36cm.
  • The full range of CVI mirror is 3.5mrad according to its datasheet and linear range of QPD is 0.2mm, so I set the distance between PZT mirrors and QPD to be about 6cm. (I realized it was wrong. When mirror tilts 1 deg, the angle of beam changes 2 deg. So the distance should be the half.)
  • After applying 0V to PZT driver input (at that time 50V was applied to PZT mirror), then centered beam spot on QPD using steering mirror, which was confirmed by monitored Pitch and Yaw signals of QPD that were around zero.  
  • In order to avoid hysteresis effect, I stared with an input signal of -10V. I then increased the input voltage in steps of 1V through the full range from -10V to +10V DC. The other input was kept 0V.
  • Both the X and Y coordinates were noted in the plot in order to investigate pitch-yaw coupling.

The calibration factor was

CVI-pitch: 0.089 mrad/V

CVI-yaw: 0.096 mrad/V

Laseroptic-pitch: 0.062 mrad/V

Laseroptic-yaw: 0.070 mrad/V

Comments:

  • I made sure that PZT mirrors move linearly in full input range (+-10V).
  • PZT CH1 input: Yaw, CH2: Pitch, CH3: +100V bias
  • The calibration factor of PZT mirrors [mrad/V] are not consistent with previous calibration (elog:40m/8967). I will check it again.
  • I measured the beam power in order to calibrate QPD responce with a powermeter, but it didn't have high precision. So I used SUM output of QPD to the calibration.
  • Full range of PZT mirrors looks 2 times smaller.

Reference:
Previous calibration of the same mirrors, elog:40/8967

Attachment 1: PZTM1calibrationCH2.pdf
PZTM1calibrationCH2.pdf
Attachment 2: PZTM1calibrationCH1.pdf
PZTM1calibrationCH1.pdf
Attachment 3: PZTM2calibrationCH2.pdf
PZTM2calibrationCH2.pdf
Attachment 4: PZTM2calibrationCH1.pdf
PZTM2calibrationCH1.pdf
  14223   Mon Oct 1 22:20:42 2018 gautamUpdateSUSPrototyping HV Bias Circuit

Summary:

I've been plugging away at Altium prototyping the high-voltage bias idea, this is meant to be a progress update.

Details:

I need to get footprints for some of the more uncommon parts (e.g. PA95) from Rich before actually laying this out on a PCB, but in the meantime, I'd like feedback on (but not restricted to) the following:

  1. The top-level diagram: this is meant to show how all this fits into the coil driver electronics chain.
    • The way I'm imagining it now, this (2U) chassis will perform the summing of the fast coil driver output to the slow bias signal using some Dsub connectors (existing slow path series resistance would simply be removed). 
    • The overall output connector (DB15) will go to the breakout board which sums in the bias voltage for the OSEM PDs and then to the satellite box.
    • The obvious flaw in summing in the two paths using a piece of conducting PCB track is that if the coil itself gets disconnected (e.g. we disconnect cable at the vacuum flange), then the full HV appears at TP3 (see pg2 of schematic). This gets divided down by the ratio of the series resistance in the fast path to slow path, but there is still the possibility of damaging the fast-path electronics. I don't know of an elegant design to protect against this.
  2. Ground loops: I asked Johannes about the Acromag DACs, and apparently they are single ended. Hopefully, because the Sorensens power Acromags, and also the eurocrates, we won't have any problems with ground loops between this unit and the fast path.
  3. High-voltage precautons: I think I've taken the necessary precautions in protecting against HV damage to the components / interfaced electronics using dual-diodes and TVSs, but someone more knowledgable should check this. Furthermore, I wonder if a Molex connector is the best way to bring in the +/- HV supply onto the board. I'd have liked to use an SHV connector but can't find a comaptible board-mountable connector.
  4.  Choice of HV OpAmp: I've chosen to stick with the PA95, but I think the PA91 has the same footprint so this shouldn't be a big deal.
  5.  Power regulation: I've adapted the power regulation scheme Rich used in D1600122 - note that the HV supply voltage doesn't undergo any regulation on the board, though there are decoupling caps close to the power pins of the PA95. Since the PA95 is inside a feedback loop, the PSRR should not be an issue, but I'll confirm with LTspice model anyways just in case.
  6. Cost: 
    • ​​Each of the metal film resistors that Rich recommended costs ~$15.
    • The voltage rating on these demand that we have 6 per channel, and if this works well, we need to make this board for 4 optics.
    • The PA95 is ~$150 each, and presumably the high voltage handling resistors and capacitors won't be cheap.
    • Steve will update about his HV supply investigations (on a secure platform, NOT the elog), but it looks like even switching supplies cost north of $1200.
    • However, as I will detail in a separate elog, my modeling suggests that among the various technical noises I've modeled so far, coil driver noise is still the largest contribution which actually seems to exceed the unsqueezed shot noise of ~ 8e-19 m/rtHz for 1W input power and PRG 40 with 20ppm RT arm losses, by a smidge (~9e-19 m/rtHz, once we take into account the fast and slow path noises, and the fact that we are not exactly Johnson noise limited).

I also don't have a good idea of what the PCB layer structure (2 layers? 3 layers? or more?) should be for this kind of circuit, I'll try and get some input from Rich.

*Updated with current noise (Attachment #2) at the output for this topology of series resistance of 25 kohm in this path. Modeling was done (in LTspice) with a noiseless 25kohm resistor, and then I included the Johnson noise contribution of the 25k in quadrature. For this choice, we are below 1pA/rtHz from this path in the band we care about. I've also tried to estimate (Attachment #3) the contribution due to (assumed flat in ASD) ripple in the HV power supply (i.e. voltage rails of the PA95) to the output current noise, seems totally negligible for any reasonable power supply spec I've seen, switching or linear.

Attachment 1: CoilDriverBias.pdf
CoilDriverBias.pdf CoilDriverBias.pdf CoilDriverBias.pdf
Attachment 2: currentNoise.pdf
currentNoise.pdf
Attachment 3: PSRR.pdf
PSRR.pdf
  14222   Mon Oct 1 20:39:09 2018 gautamConfigurationASCc1asy

We need to set up a copy of the c1asx model (which currently runs on c1iscex), to be named c1asy, on c1iscey for the green steering PZTs. The plan discussed at the meeting last Wednesday was to rename the existing model c1tst into c1asy, and recompile it with the relevant parts copied over from c1asx. However, I suspect this will create some problems related to the "dcuid" field in the CDS params block (I ran into this issue when I tried to use the dcuid for an old model which no longer exists, called c1imc, for the c1omc model).

From what I can gather, we should be able to circumvent this problem by deleting the .par file corresponding to the c1tst model living at /opt/rtcds/caltech/c1/target/gds/param/, and rename the model to c1asy, and recompile it. But I thought I should post this here checking if anyone knows of other potential conflicts that will need to be managed before I start poking around and breaking things. Alternatively, there are plenty of cores available on c1iscey, so we could just set up a fresh c1asy model...

 
  • (write programming code of making alignment control automatically)
  14221   Mon Oct 1 13:33:55 2018 yukiConfigurationASCQPD calibration
Quote:

I assume this QPD set is a D1600079/D1600273 combo.

How much was the SUM output during the measurement? Also how much were the beam radii of this beam (from the error func fittings)?
Then the calibration [V/m] is going to be the linear/inv-linear function of the incident power and the beam radus.

You mean the linear range is +/-50mV (for a given beam), I guess.

  • The SUM output was from -174 to -127 mV.
  • The beam radii calculated from the error func fittings was 0.47 mm.
  • Total optical path length measured by a ruler= 36 cm.
  • Beam power measured at QPD was 2.96 mW. (There are some loss mechanism in the setup.)

Then the calibration factor of the QPD is

X axis: 584 * (POWER / 2.96mW) * (0.472mm /  RADIUS) [mV/mm]
Y axis: 588 * (POWER / 2.96mW) * (0.472mm /  RADIUS) [mV/mm].

Attachment 1: Pic_QPDcalibration.jpg
Pic_QPDcalibration.jpg
  14220   Mon Oct 1 12:03:41 2018 not yukiConfigurationASCPZT driver board verification

I assume this QPD set is a D1600079/D1600273 combo.

How much was the SUM output during the measurement? Also how much were the beam radii of this beam (from the error func fittings)?
Then the calibration [V/m] is going to be the linear/inv-linear function of the incident power and the beam radus.

You mean the linear range is +/-50mV (for a given beam), I guess.

 

  14219   Sun Sep 30 20:14:51 2018 yukiConfigurationASCQPD calibration

[ Yuki, Gautam, Steve ]

Results:
I calibrated a QPD (D1600079, V1009) and made sure it performes well. The calibration constants are as follows:

X-Axis: 584 mV/mm
Y-Axis: 588 mV/mm

Details:
The calibration of QPD is needed to calibrate steeing PZT mirrors. It was measured by moving QPD on a translation stage. The QPD was connected to its amplifier (D1700110-v1) and +-18V was supplied from DC power supplier. The amplifier has three output ports; Pitch, Yaw, and Sum. I did the calibration as follows:

  • Center beam spot on QPD using steering mirror, which was confirmed by monitored Pitch and Yaw signals that were around zero.  
  • Kept Y-axis micrometer fixed, moved X-axis micrometer and measured the outputs. 
  • Repeated the procedure for the Y-axis. 

The results are attached. The main signal was fitted with error function and I drawed a slope at zero crossing point, which is calibration factor. I determined the linear range of the QPD to be when the output was in range -50V to 50V, then corresponding displacement range is about 0.2 mm width. Using this result, the PZT mirrors will be calibrated in linear range of the QPD tomorrow. 

Comments:

  • Some X-Y coupling existed. When one axis micrometer was moved, a little signal of the other direction was also generated.
  • As Gautam proposed in the previous study, there is some hysteresis. That process would bring some errors to this result.
  • A scale of micrometer is expressed in INCH!
  • The micrometer I used was made to have 1/2 inch range, but it didn't work well and the range of X-axis was much narrower. 

Reference:
previous experiment by Gautam for X-arm: elog:40m/8873, elog:40m/8884

Attachment 1: QPDcalibrationXaxis.pdf
QPDcalibrationXaxis.pdf
Attachment 2: QPDcalibrationYaxis.pdf
QPDcalibrationYaxis.pdf
  14218   Thu Sep 27 14:02:55 2018 yukiConfigurationASCPZT driver board verification

[ Yuki, Gautam ]

I fixed the input terminal that had been off, and made sure PZT driver board performs as we expect. 

At first I ran a simulation of the PZT driver circuit using LTspice (Attached #1 and #2). It shows that when the bias is 30V the driver performs well only with high input volatage (bigger than 3V). Then I measured the performance as following way:

  1. Applied +-15V to the board with an expansion card and 31.8V to the high voltage port which is the maximum voltage of PS280 DC power supplier C10013.
  2. Terminated input and connectd input bias to GND, then set offset to -10.4V. This value is refered as elog:40m/8832.
  3. Injected DC signal into input port using a function generator.
  4. Measured voltage at the OUT port and MON port.

The result of this is attached #3 and #4. It is consistent with simulated one. All ports performed well.

  • V(M1_PIT_OUT) = -4.86 *Vin +49.3 [V]
  • V(M1_YAW_OUT) = -4.86 *Vin +49.2 [V]
  • V(M2_PIT_OUT) = -4.85 *Vin +49.4 [V]
  • V(M2_YAW_OUT) = -4.86 *Vin +49.1 [V]
  • V(M1_PIT_MON) = -0.333 *Vin +3.40 [V]
  • V(M1_YAW_MON) = -0.333 *Vin +3.40 [V]
  • V(M2_PIT_MON) = -0.333 *Vin +3.40 [V]
  • V(M2_YAW_MON) = -0.333 *Vin +3.40 [V]

The high voltage points (100V DC) remain to be tested.

Attachment 1: PZTdriverSimulationDiagram.pdf
PZTdriverSimulationDiagram.pdf
Attachment 2: PZTdriverSimulationResult.pdf
PZTdriverSimulationResult.pdf
Attachment 3: PZTdriverPerformanceCheck_ResultOUT.pdf
PZTdriverPerformanceCheck_ResultOUT.pdf
Attachment 4: PZTdriverPerformanceCheck_ResultMON.pdf
PZTdriverPerformanceCheck_ResultMON.pdf
Attachment 5: PZTdriver.asc
Version 4
SHEET 1 2120 2120
WIRE 1408 656 1408 624
WIRE 1552 656 1552 624
WIRE 1712 656 1712 624
WIRE 1872 656 1872 624
WIRE 2016 656 2016 624
WIRE 1408 768 1408 736
WIRE 1552 768 1552 736
WIRE 1712 768 1712 736
... 193 more lines ...
ELOG V3.1.3-